# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 7.1 Again scalar condensate
where parameters $c_{1}$, $c_{2}$ and $c_{3}$ are expressed through values of the binding energy $\mathcal{E}_{0}=-\mathcal{E}_{N\rm}(n_{0})$, compressibility modulus $K$ of the nuclear matter and the nuclear saturation density $n_{0}$. Taking $m_{N}=939$ MeV for the free nucleon mass and appropriate values $n_{0}=0.17\,{\rm fm}^{-3}$, $\mathcal{E}_{0}=-16$ MeV and $K=285$ MeV we find Kolomeitsev and Voskresensky (2016), $c_{1}=3.946$ , $c_{2}=3.837$$c_{3}=13.10$ . From Eq. (45) the volume part of the nucleon contribution to the scalar Landau-Migdal parameter $f_{0}^{N}\equiv f_{0}$ follows as
$f_{0}(n)=\frac{3m_{N}^{*}}{p_{\rm F}^{2}}n\frac{{\rm d}^{2}(n\mathcal{E}_{N,\rm v})}{{\rm d}n^{2}}-1=\frac{m_{N}^{*}}{m_{N}}-1+\frac{m_{N}^{*}}{m_{N}}\frac{p_{\rm F}}{m_{N}}\Big(-6c_{1}+c_{2}\frac{28p_{\rm F}}{3m_{N}}+c_{3}\frac{40p_{\rm F}^{2}}{3m_{N}^{2}}\Big)\,.$ (46)
The energy per particle and the scalar Landau-Migdal parameter are shown in Fig. 8 as functions of the nucleon density. The nucleon volume part of the energy per particle and the parameter $f_{0}$ given by Eq. (46) are presented on the left panel in Fig. 8 by the dotted lines. There is a broad interval of densities, $n_{c}^{\rm(l)}<n<n_{c}^{\rm(u)}$ where $f_{0}<-1$ with $n_{c}^{\rm(l)}=0.422\times 10^{-2}\,n_{0}$ and $n_{c}^{\rm(u)}=0.655\,n_{0}$. The condensate contribution proves to be strongly attractive creating a minimum at $n\sim 0.4\,n_{0}$ for a reduced value of $\lambda$ for $\xi<1/2$. The energies per nucleon $\mathcal{E}_{N}+\mathcal{E}_{B}$ are shown on the left upper panel by thin lines labeled by values of the parameter $\xi$. They are calculated according to Eq. (44) with $f_{0}(n)$ given by Eq. (46).
@@
We used the variational approach to find the solution of Eq. (47), see Ref. Kolomeitsev and Voskresensky (2016), and we fixed the values of $f_{0}^{\rm tot}$ at the ends of the instability interval $f_{0}^{\rm tot}(n_{c}^{(l)})=f_{0}^{\rm tot}(n_{c}^{(u)})=-1$. Thereby we postulated that the condensate arises by the second-order phase transition. Taking into account the possibility of a first-order phase transition, the energy could be further decreased. The energy per particle in our case is constructed using Eq. (32):
@@
where we used $\tilde{\omega}^{2}_{s}(0)=-1-1/f_{0}^{\rm tot}$. Thus, if originally $f_{0}<-1$ and therefore $f_{0}^{\rm tot}<-1$, the renormalized interaction yields $1<f_{\rm 0,ren}<-1/2$. Hence, the first-sound modes becomes stable in the Fermi liquid with the condensate. The quantity $f_{\rm 0,ren}$ is plotted on the right lower panel in Fig. 8. Knowing the value $f_{\rm 0,ren}$ one can reconstruct the energy per particle $\mathcal{E}_{\rm tot}(n)$ of the system using Eq. (48) after the replacement $f_{0}^{\rm tot}\to f_{\rm 0,ren}$ in the integrand. The results are shown on the right upper panel in Fig. 8. We observe the appearance a minimum of the function $\mathcal{E}_{\rm tot}(n)$. For $\xi=1$ the minimum is at $n_{\rm min}=0.085\,n_{0}$ with $\mathcal{E}_{\rm min}=-16.01$ MeV. Variation of the parameter $\xi$ in interval $\xi=1/4\mbox{--}2.0$ leads to variations of the values $n_{\rm min}=(0.82\mbox{--}0.93)\times 10^{-1}\,n_{0}$ with $\mathcal{E}_{\rm min}=-(16.8\mbox{--}15.6)$ MeV. So the minimum in energy can correspond to a new stable or metastable state of the dilute nuclear matter compared with the usual saturation energy of 16-16 MeV used to specify the equation of state of the ISM.
@@
Considering finite system one must take into account the surface and Coulomb energies. We parameterize the surface energy as in Migdal et al. (1976); Migdal (1978),
$\mathcal{E}_{\rm S}A=a_{\rm S}A^{2/3}\,,\quad a_{\rm S}\approx C_{\rm S}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{\mathcal{E}(n,z)}{\mathcal{E}(n_{0},z=1/2)}\,,\quad C_{\rm S}=18\,{\rm MeV}\,$, (51)
@@
$f_{0}^{\rm S}(n)$ $=f_{0}(n)+\frac{C_{\rm S}}{A^{1/3}}\frac{1}{\mathcal{E}_{N,\rm v}(n_{0})}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{m_{N}^{*}}{m_{N}}\Big(\frac{3}{5}-c_{1}\frac{28p_{\rm F}}{9m_{N}}+c_{2}\frac{40p_{\rm F}^{2}}{9m_{N}^{2}}+c_{3}\frac{6p_{\rm F}^{3}}{m_{N}^{3}}\Big)\,.$ (52)
Following  Kolomeitsev and Voskresensky (2016), as an example, we consider a system of 125 nucleons at $N\approx Z$. Then the energy per particle $\mathcal{E}_{N}+\mathcal{E}_{\rm S}$ saturates at density $0.95\,n_{0}$ with the minimum $\mathcal{E}_{0}^{\rm S}=-12.4$ MeV. With taking into account the surface energy and solving Eq. (47) with $f_{0}^{\rm S}$ instead of $f_{0}$ we find that the new minimum appears at $n_{\rm min}=0.74\times 10^{-1}\,n_{0}$ with the depth $\mathcal{E}_{\rm min}=-13.04$ MeV. So, $\mathcal{E}_{\rm min}<\mathcal{E}_{0}^{\rm S}$ and, hence, a new stable state of matter is possible in the finite size object with account for the surface energy. To analyze its stability against fusion we must include the Coulomb energy
$\mathcal{E}_{\rm C}A\simeq a_{\rm C}z^{2}A^{5/3}\,,\quad a_{\rm C}=C_{\rm C}\Big(\frac{n}{n_{0}}\Big)^{1/3}\,,\quad C_{\rm C}=0.7\,{\rm MeV}\,.$ (53)
@@
We see that, because of the condensate formation, a new metastable dilute nuclear state is stable against spontaneous the fission for $A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}170$.
Reference Maslov and Voskresensky (2019) studied the equation of state of isospin-asymmetric matter and demonstrated that even for the asymmetry parameter $n_{p}/n_{n}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.2$ there remains a broad region of proton and neutron densities where $f_{0}<-1$ and, therefore, the spinodal instability may develop. For a nugget of a size $R$ much larger than the Debye screening length $l_{D}\simeq\sqrt{\pi}/(2|e|V_{0})$, the Coulomb contribution from protons must be screened by electrons, $n_{p}=n_{e}$, cf. Migdal et al. (1977). Here we for simplicity assume that the scalar field is electrically neutral. The electron Fermi momentum is then $p_{{\rm F},e}=V_{0}=(3\pi^{2}n_{p})^{1/3}$ and the electron kinetic energy density is $E_{e}=p_{{\rm F},e}^{4}/(4\pi^{2})$. In the isospin-asymmetric matter, the nuclear symmetry energy must be included. A simple parametrization was proposed in Heiselberg and Hjorth-Jensen (1999), which is a fit to microscopic calculations
$E_{N,{\rm sym}}(n_{p},n_{n})=J\,\frac{(n_{p}+n_{n})^{\gamma}}{n_{0}^{\gamma}}\frac{(n_{p}-n_{n})^{2}}{n_{p}+n_{n}}\,,\quad J=32\,{\rm MeV}\,,\quad\gamma=0.6\,.$ (55)
@@
$\mathcal{E}(\chi)$ $=\mathcal{E}_{\rm min}+J\,\Big(\frac{n_{\rm min}}{n_{0}}\Big)^{\gamma}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}+\frac{3}{4}(3\pi^{2}n_{\rm min})^{1/3}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}$
$\approx\mathcal{E}_{\rm min}+7.3\,{\rm MeV}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}+149\,{\rm MeV}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}\,$, (56)
where $n_{\rm min}=0.85\times 10^{-1}\,n_{0}$ is used. If we take $\chi\sim 0.2$ we obtain $\mathcal{E}(\chi\sim 0.2)\approx\mathcal{E}_{\rm min}+16.9\,{\rm MeV}$. Minimizing $\mathcal{E}(\chi)$ with respect to $\chi$ we find that the minimum is realized for $\chi_{\rm min}\approx 3.1\times 10^{-3}$ and in the minimum $\mathcal{E}(\chi_{\rm min})\approx\mathcal{E}_{\rm min}+7.3\,{\rm MeV}$. Thus if $\mathcal{E}_{\rm min}$ is of the order (1017)-(10-17) MeV in the case of the isospin asymmetric nuclear matter, one may rise the question about a possibility of the existence of neutron-rich dilute compact stars glued by an collective scalar field. However, the above estimates crucially depend on the value of symmetry energy of nuclear matter at small densities: the smaller $E_{N,{\rm sym}}$ is, the smaller will be repulsive contributions in Eq. (56).
@@
7.2 Bose condensates of nuclear clusters
At densities when $f_{0}(n)<-1$ the incompressibility of nuclear matter is negative and, as we have mentioned, not only to the scalar particle-hole mode condensation but also the spinodal instability of matter may occur. In the low energy heavy-ion collisions this instability manifests itself in an abrupt increase of the production of $\alpha$ particles and other nuclear clusters, see Röpke et al. (1982); Curtin et al. (1983); Bertsch and Siemens (1983); Schulz et al. (1983); Maslov and Voskresensky (2019); Borderie et al. (2022). With a binding energy of 7.07 MeV per nucleon $\alpha$ particles are bound as strongly as the most tightly bound nucleus of 52Fe. Under certain conditions $\alpha$ particles can form a stable Bose-Einstein condensate state (see Ref. Typel et al. (2010); Schuck et al. (2016) and references therein). Reference Clark and Krotscheck (2024) argued that the spectrum of excitations in the system of interacting $\alpha$ particles is similar to the spectrum of He4 and has the roton minimum at $k\neq 0$. If so, the inhomogeneous $\alpha$ condensate occupying the roton minimum could occur in the moving (e.g. rotating) pieces of the $\alpha$ particle system. Similar effects were studied in Refs. Voskresensky (1993, 2023); Kolomeitsev and Voskresensky (2017). Also, one considers the possibility of deuteron condensation in dilute nuclear systems as discussed in Ref. Lombardo et al. (2001), despite the fact that deuterons are seven times weakly bound than $\alpha$ particles.
Consider an extended nucleon system of large size and study whether it is possible to form a Bose-Einstein condensate from bosonic clusters consisting of an even number of nucleons. We must take into account an electric field and $\beta$ equilibrium conditions, as it was done in Ref. Voskresensky (1977); Migdal et al. (1990) and discussed above in Section 3.
@@ 7.2 Bose condensates of nuclear clusters
where $\omega+V$ is the energy of a positively charged cluster quasiparticle shifted by the electric potential $V$. Considering a large system, we may drop gradient terms with $\nabla\phi$ and $\nabla V$ and neglect for $\phi$ self-interaction terms for simplicity. The equation of motion for the scalar field following from the Lagrangian (57) is $(\omega+V)^{2}\phi=m^{2}_{cl}\phi$. The condition of the charge neutrality of the system requires that the positive charge of clusters is compensated by the charge of electrons, therefore $n_{\rm cl}=n_{e}=V^{3}/(3\pi^{2})$, where the number density of cluster quasiparticles is
@@
$\Delta E=E_{\rm tot}-A_{\rm cl}m_{N}n_{\rm cl}\simeq\frac{V^{4}}{12\pi^{2}}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}n_{\rm cl}=\Big(\frac{1}{4}(3\pi^{2}n_{\rm cl})^{1/3}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}\Big)n_{\rm cl}\,.$ (58)
@@
Thus, in this simplified model, the system is unbound ($\Delta E>0$) at densities,
$n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(4A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)})^{3}/(3\pi^{2})$
@@
in the large system limit. For $\alpha$ particles it occurs at densities $n>0.04\,n_{0}$.
7.3 Dilute nonequilibrium states

### v2

@@ VII.1 Again, a scalar condensate
where parameters $c_{1}$, $c_{2}$ and $c_{3}$ are expressed through values of the binding energy $\mathcal{E}_{0}=-\mathcal{E}_{N\rm}(n_{0})$, compressibility modulus $K$ of the nuclear matter and the nuclear saturation density $n_{0}$. Taking $m_{N}=939$ MeV for the free nucleon mass and appropriate values $n_{0}=0.17\,{\rm fm}^{-3}$, $\mathcal{E}_{0}=-16$ MeV and $K=285$ MeV we find [158], $c_{1}=3.946$ , $c_{2}=3.837$$c_{3}=13.10$ . From Eq. (56) the volume part of the nucleon contribution to the scalar Landau-Migdal parameter $f_{0}$ follows as
$f_{0}(n)$ $=\frac{3m_{N}^{*}}{p_{\rm F}^{2}}n\frac{{\rm d}^{2}(n\mathcal{E}_{N,\rm v})}{{\rm d}n^{2}}-1=\frac{m_{N}^{*}}{m_{N}}-1$
$+\frac{m_{N}^{*}}{m_{N}}\frac{p_{\rm F}}{m_{N}}\Big(-6c_{1}+c_{2}\frac{28p_{\rm F}}{3m_{N}}+c_{3}\frac{40p_{\rm F}^{2}}{3m_{N}^{2}}\Big)\,.$ (57)
The energy per particle and the scalar Landau-Migdal parameter are shown in Fig. 11 as functions of the nucleon density. The nucleon volume part of the energy per particle and the parameter $f_{0}$ given by Eq. (57) are presented on the left panel in Fig. 11 by the dotted lines. There is a broad interval of densities, $n_{c}^{\rm(l)}<n<n_{c}^{\rm(u)}$ where $f_{0}<-1$ with $n_{c}^{\rm(l)}=0.422\times 10^{-2}\,n_{0}$ and $n_{c}^{\rm(u)}=0.655\,n_{0}$. Similar dependence $f_{0}(n)$ is obtained for the DD2 equation of state, see Fig. 3 in [239]. Notice that taking into account of the clustering effects [237, 238] may affect this dependence, see Fig. 9 in [239]. We will focus on the alternative taking into account effect of possible Bose condensation. The condensate contribution proves to be strongly attractive creating a minimum at $n\sim 0.4\,n_{0}$ for a reduced value of $\lambda$ for $\xi<1/2$. The energies per nucleon $\mathcal{E}_{N}+\mathcal{E}_{B}$ are shown on the left upper panel by thin lines labeled by values of the parameter $\xi$. They are calculated according to Eq. (55) with $f_{0}(n)$ given by Eq. (57).
@@
We used the variational approach to find the solution of Eq. (58), see Ref. [158], and we fixed the values of $f_{0}^{\rm tot}$ at the ends of the instability interval $f_{0}^{\rm tot}(n_{c}^{(l)})=f_{0}^{\rm tot}(n_{c}^{(u)})=-1$. Thereby we postulated that the condensate arises by the second-order phase transition. Taking into account the possibility of a first-order phase transition, the energy could be further decreased. The energy per particle in our case is constructed using Eq. (43):
@@
where we used $\tilde{\omega}^{2}_{s}(0)=-1-1/f_{0}^{\rm tot}$. Thus, if originally $f_{0}<-1$ and therefore $f_{0}^{\rm tot}<-1$, the renormalized interaction yields $1<f_{\rm 0,ren}<-1/2$. Hence, the first-sound modes become stable in the Fermi liquid with the condensate. The quantity $f_{\rm 0,ren}$ is plotted on the right lower panel in Fig. 11. Knowing the value $f_{\rm 0,ren}$ one can reconstruct the energy per particle $\mathcal{E}_{\rm tot}(n)$ of the system using Eq. (59) after the replacement $f_{0}^{\rm tot}\to f_{\rm 0,ren}$ in the integrand. The results are shown on the right upper panel in Fig. 11. We observe the appearance of a minimum of the function $\mathcal{E}_{\rm tot}(n)$. For $\xi=1$ the minimum is at $n_{\rm min}=0.085\,n_{0}$ with $\mathcal{E}_{\rm min}=-16.01$ MeV. Variation of the parameter $\xi$ in interval $\xi=1/4\mbox{--}2.0$ leads to variations of the values $n_{\rm min}=(0.82\mbox{--}0.93)\times 10^{-1}\,n_{0}$ with $\mathcal{E}_{\rm min}=-(16.8\mbox{--}15.6)$ MeV. So the minimum in energy can correspond to a new stable or metastable state of the dilute nuclear matter compared with the usual saturation energy of 16-16 MeV used to specify the equation of state of the ISM.
@@
Considering finite system one must take into account the surface and Coulomb energies. We parameterize the surface energy as in [211, 215],
$\mathcal{E}_{\rm S}A$ $=a_{\rm S}A^{2/3}\,,\quad a_{\rm S}\approx C_{\rm S}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{\mathcal{E}(n,z)}{\mathcal{E}(n_{0},z=1/2)}\,$$\quad$
$C_{\rm S}$ $=18\,{\rm MeV}\,$, (62)
@@
$f_{0}^{\rm S}(n)$ $=f_{0}(n)+\frac{C_{\rm S}}{A^{1/3}}\frac{1}{\mathcal{E}_{N,\rm v}(n_{0})}\Big(\frac{n}{n_{0}}\Big)^{1/3}\frac{m_{N}^{*}}{m_{N}}$
$\times\Big(\frac{3}{5}-c_{1}\frac{28p_{\rm F}}{9m_{N}}+c_{2}\frac{40p_{\rm F}^{2}}{9m_{N}^{2}}+c_{3}\frac{6p_{\rm F}^{3}}{m_{N}^{3}}\Big)\,.$ (63)
Following  [158], as an example, we consider a system of 125 nucleons at $N\approx Z$. Then the energy per particle $\mathcal{E}_{N}+\mathcal{E}_{\rm S}$ saturates at density $0.95\,n_{0}$ with the minimum $\mathcal{E}_{0}^{\rm S}=-12.4$ MeV. With taking into account the surface energy and solving Eq. (58) with $f_{0}^{\rm S}$ instead of $f_{0}$ we find that the new minimum appears at $n_{\rm min}=0.74\times 10^{-1}\,n_{0}$ with the depth $\mathcal{E}_{\rm min}=-13.04$ MeV. So, $\mathcal{E}_{\rm min}<\mathcal{E}_{0}^{\rm S}$ and, hence, a new stable state of matter is possible in the finite size object with account for the surface energy. To analyze its stability against fusion we must include the Coulomb energy
$\mathcal{E}_{\rm C}A\simeq a_{\rm C}z^{2}A^{5/3}\,,\,\,a_{\rm C}=C_{\rm C}\Big(\frac{n}{n_{0}}\Big)^{1/3}\,,\,\,C_{\rm C}=0.7\,{\rm MeV}\,.$ (64)
@@
We see that, because of the condensate formation, a new metastable dilute nuclear state is stable against spontaneous fission for $A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}170$.
Reference [198] studied the equation of state of isospin-asymmetric matter and demonstrated that even for the asymmetry parameter $n_{p}/n_{n}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.2$ there remains a broad region of proton and neutron densities where $f_{0}<-1$ and, therefore, the spinodal instability may develop. For a nugget of a size $R$ much larger than the Debye screening length $l_{D}\simeq\sqrt{\pi}/(2|e|V_{0})$, the Coulomb contribution from protons must be screened by electrons, $n_{p}=n_{e}$, cf. [209]. Here we for simplicity assume that the scalar field is electrically neutral. The electron Fermi momentum is then $p_{{\rm F},e}=V_{0}=(3\pi^{2}n_{p})^{1/3}$ and the electron kinetic energy density is $E_{e}=p_{{\rm F},e}^{4}/(4\pi^{2})$. In the isospin-asymmetric matter, the nuclear symmetry energy must be included. A simple parametrization was proposed in [117], which is a fit to microscopic calculations
$E_{N,{\rm sym}}(n_{p},n_{n})$ $=J\,\frac{(n_{p}+n_{n})^{\gamma}}{n_{0}^{\gamma}}\frac{(n_{p}-n_{n})^{2}}{n_{p}+n_{n}}\,$$\quad$
$J$ $=32\,{\rm MeV}\,,\quad\gamma=0.6\,.$ (66)
@@
$\mathcal{E}(\chi)=\mathcal{E}_{\rm min}+J\,\Big(\frac{n_{\rm min}}{n_{0}}\Big)^{\gamma}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}$
$+\frac{3}{4}(3\pi^{2}n_{\rm min})^{1/3}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}\approx\mathcal{E}_{\rm min}$
$+7.3\,{\rm MeV}\frac{(1-\chi)^{2}}{(1+\chi)^{2}}+149\,{\rm MeV}\frac{\chi^{4/3}}{(1+\chi)^{4/3}}\,$, (67)
where $n_{\rm min}=0.85\times 10^{-1}\,n_{0}$ is used. If we take $\chi\sim 0.2$ we obtain $\mathcal{E}(\chi\sim 0.2)\approx\mathcal{E}_{\rm min}+16.9\,{\rm MeV}$. Minimizing $\mathcal{E}(\chi)$ with respect to $\chi$ we find that the minimum is realized for $\chi_{\rm min}\approx 3.1\times 10^{-3}$ and in the minimum $\mathcal{E}(\chi_{\rm min})\approx\mathcal{E}_{\rm min}+7.3\,{\rm MeV}$. Thus if $\mathcal{E}_{\rm min}$ is of the order (1017)-(10-17) MeV in the case of the isospin asymmetric nuclear matter, one may rise the question about a possibility of the existence of neutron-rich dilute compact stars glued by an collective scalar field. However, the above estimates crucially depend on the value of symmetry energy of nuclear matter at small densities. The smaller $E_{N,{\rm sym}}$ is, the smaller will be repulsive contributions in Eq. (67).
@@
VII.2 Bose condensates of nuclear clusters
At densities when $f_{0}(n)<-1$ the incompressibility of nuclear matter is negative and, as we have mentioned, not only the scalar particle-hole mode condensati on may occur but also the spinodal instability of matter. In the low energy heavy-ion collisions the latter instability manifests itself in an abrupt increase of the production of $\alpha$ particles and other nuclear clusters, see [235, 60, 27, 254, 198, 39]. With a binding energy of 7.07 MeV per nucleon $\alpha$ particles are bound as strongly as the most tightly bound nucleus of 52Fe. Under certain conditions $\alpha$ particles can form a stable Bose-Einstein condensate state (see Ref. [270, 236, 253] and references therein). Reference [54] argued that the spectrum of excitations in the system of interacting $\alpha$ particles is similar to the spectrum of He4 and has the roton minimum at $k\neq 0$. If so, the inhomogeneous $\alpha$ condensate occupying the roton minimum could occur in the moving (e.g. rotating) pieces of the $\alpha$ particle system. Similar effects were studied in Refs. [283, 288, 160], see discussion in Sect. III.3. Also, one considers the possibility of deuteron condensation in dilute nuclear systems as discussed in Ref. [188], despite the fact that deuterons are seven times weakly bound than $\alpha$ particles.
Consider an extended nucleon system of large size and study whether it is possible to form a Bose-Einstein condensate from bosonic clusters consisting of an even number of nucleons. We must take into account an electric field and $\beta$ equilibrium conditions, as it was done in Ref. [282, 210] and discussed above in Section III.
@@ VII.2 Bose condensates of nuclear clusters
where $\omega+V$ is the energy of a positively charged cluster quasiparticle shifted by the electric potential $V$. Here, considering a large system, we dropped gradient terms with $\nabla\phi$ and $\nabla V$ and neglected $\phi$ self-interaction terms for simplicity, similarly to that we have done finding solutions in Sects. III.1 and III.3. The equation of motion for the scalar field following from the Lagrangian (68) is $(\omega+V)^{2}\phi=m^{2}_{cl}\phi$. The condition of the charge neutrality of the system requires that the positive charge of clusters is compensated by the charge of electrons, therefore $n_{\rm cl}=n_{e}=V^{3}/(3\pi^{2})$, where the number density of cluster quasiparticles is
@@
$\Delta E$ $=E_{\rm tot}-A_{\rm cl}m_{N}n_{\rm cl}\simeq\frac{V^{4}}{12\pi^{2}}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}n_{\rm cl}$
$=\Big(\frac{1}{4}(3\pi^{2}n_{\rm cl})^{1/3}-A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)}\Big)n_{\rm cl}\,.$ (69)
@@
Thus, in this simplified model, the system is bound ($\Delta E<0$) at densities
$n<n_{\rm Mott}=(4A_{\rm cl}\mathcal{E}_{\rm cl}^{\rm(bind)})^{3}/(3\pi^{2})$
@@
in the large-system limit, and it is unbound for $n>n_{\rm Mott}$. For $\alpha$ particles we find $n_{\rm Mott}\simeq 0.04\,n_{0}$.
VII.3 Dilute nonequilibrium “anomalous” states